OpenStudy (anonymous):

Some birds and squirrels were on a branch of a tree. There were 15 animals on the branch with a total of 50 animal feet. How many birds and how many squirrels were on the branch? Part 1: Set up the problem. Part 2: Explain why you set it up that way. Part 3: Provide the solution.

OpenStudy (anonymous):

donn't mind me lagrange. I'm just here for the relaxing atmosphere.

OpenStudy (anonymous):

simple system of equations: in total ofcourse there are 50 animal feet so: x+y=50 (where x and y are squirrels and birds respectively) Then second equation would be: Keeping in mind that in this world, squirrels have 4 feet and birds have 2 feet is: 4x+2y=15. So the system ends up being: {x+y=50} {4x+2y=15} Now can you solve this?

OpenStudy (anonymous):

Yeah, Thank you so much. (:

OpenStudy (anonymous):

just do it quietly ... don't disturb the wildlife ...

OpenStudy (anonymous):

Typo: The first equation should be equal to 15 and the second should be equal to 50: {x+y=15} {4x+2y=50}